# Scoring Nim

### v1

@@ 6 Changes in the Optimal Strategy Depending on the Bonus Value
Thus, it follows that $f_{N}(2m+1,2m,1)\geq 2$ for $N=-(2m-2),-(2m-4),\ldots,2m-4,2m-2$. Furthermore, since the slope of this function with respect to $N$ is at least $-1$ (as stated in Proposition 3.1(iii)), it follows that $f_{N}(2m+1,2m,1)\geq F_{m}(N)$ holds for all $N$.
Next, when $N=\pm(2l+1)$ and $10\leq l<m-$ 1, we analyze the possible moves available to the first player at the position $(2m+1,2m,1)$ and explicitly describe the opponent’s responses. Through this analysis, we show that $f_{\pm(2l+1)}(2m+1,2m,1)$ does not exceed 1.
  • If the first player moves to $(2m+1,2m,0)$, the second player will respond with a move to $(2m,2m,0)$. Then, by (3), it follows that
    $1-f_{\pm(2l+1)}(2m+1,2m,0)\leq f_{\pm(2l+1)}(2m,2m,0)=1-|1+N|\leq 1$
    Similarly, if the first player moves to $(2m,2m,1)$, the second player will again move to $(2m,2m,0)$.
  • If the first player moves to $(0,2m,1)$, the second player will respond with a move to $(0,1,1)$ if $N=2l+1$, leading to
    $2m+1-f_{2l+1}(0,2m,1)\leq 2+f_{2l+1}(0,1,1)=2-N.$
    If $N=-2l-1$, the second player will respond with a move to $(0,0,1)$, leading to
    $2m+1-f_{-2l-1}(0,2m,1)\leq 1+f_{-2l-1}(0,0,1)=2+N.$
    Thus, in both cases, the function value does not exceed 1. Similarly, if the first player moves to $(1,2m,1),(2m+1,0,1)$ or $(2m+1,1,1)$, the second player will respond with a move to $(0,1,1)$ if $N=2l+1$, and to either $(1,1,1)$ or $(0,0,1)$ if $N=-2l-1$, so that the function value does not exceed 1.
  • If the first player moves to $(x,2m,1)$ or $(2m+1,x,1)$ (where $x$ is an even number with $2\leq x\leq 2l$), the second player will respond with a move to $(x,x+1,1)$ (in this sequence of moves, it is more advantageous to have initially moved to $(x,2m,1)$ for the first player), leading to
@@
    By the induction hypothesis, i.e., by Corollary 6.4(i), we have
    $f_{\pm(2l+1)}(x,x+1,1)\leq-1$, implying that the right-hand side of the inequality does not exceed 1.
  • If the first player moves to $(x,2m,1)$ or $(2m+1,x,1)$ (where $x$ is an even number with $22l+2\leq x\leq 2m-$ 2), the second player will respond with a move to $(x,2l,1)$ (it is more advantageous to have initially moved to $(x,2m,1)$ for the first player), leading to
@@
    By the induction hypothesis, i.e., by Corollary 6.4(ii), we have
    $f_{\pm 2l}(x,2l,1)=x-2l-1$, and thus we obtain $f_{\pm(2l+1)}(x,2l,1)\leq x-2l$. Thus, the right-hand side of the above inequality does not exceed 1.
  • If the first player moves to $(x,2m,1)$ or $(2m+1,x,1)$ (where $x$ is an odd number with $13\leq x\leq 2m-$ 1), the second player will respond with a move to $(x,x-1,1)$ (it is more advantageous to have initially moved to $(x,2m,1)$ for the first player), leading to
    $2m+1-x-f_{\pm(2l+1)}(x,2m,1)\leq f_{\pm(2l+1)}(x,x-1,1).$
    By the induction hypothesis, i.e., by Corollary 6.4(i), it follows that
    the right-hand side of the above inequality does not exceed 1.
From the above moves and responses, we have established that $f_{N}(2m+1,2m,1)\leq 1$ for $N=-(2m-3),-(2m-5),\ldots,2m-5,2m-3$. Consequently, it follows that $f_{N}(2m+1,2m,1)\leq F_{m}(N)$ for $|N|\leq 2m-2$. On the other hand, by (5), it is known that for sufficiently large $|N|$, we have $f_{N}(2m+1,2m,1)=2m-|N|$. Thus, by Corollary 3.2, we conclude that for all $N$, $f_{N}(2m+1,2m,1)\leq 2m-|N|$. Hence, for all $N$, it follows that $f_{N}(2m+1,2m,1)\leq F_{m}(N)$.
Similarly, (ii) and (iii) for $k=m$ follows by assuming that (i), (ii), and (iii) hold for all $k<m$ (details omitted).  $\Box$
We provide additional insights into Theorem 6.3 and examine the factors contributing to the complexity of move selection and the increase in the number of breakpoints. First, as observed in the proof of (i), in the position $(2k+1,2k,1)$ ($k\geq 1$), the best move (or one of the best moves) for the player is to choose the even number $j$ which is closest to $N$ within the range $-(2k-2)\leq j\leq 2k-2$ and move to $(|j|,2k,1)$. For example, in the position $(9,8,1)$, the player should make the following moves based on the value of $|N|$:
@@
This strategy is derived from Corollary 6.4(ii), which states that $f_{N}(x,|j|,1)$ attains its minimum value $x-|j|-1$ at $N=\pm|j|$. This represents the weakest positions for the current player. Thus, it is advantageous to force the opponent into such a position.
Additionally, in the positions $(2k+1,2k,1)$ and $(x,2k,1)$ ($x\geq 2k+2$), considered in (i) and (ii), a common strategy is observed when $|N|$ is small: the player takes stones from the largest pile so that the remaining number of stones is close to $|N|$. As a result, the function $F_{k}(N)$, with an upward or downward shift, appears in both payoff functions. The distinction between these two payoff functions lies in whether the player can force the opponent into a P-position when $|N|$ is large. Specifically, when $|N|\geq 2k$, in the position $(2k+1,2k,1)$, the optimal move is to $(2k-2,2k,1)$, which is the same as in the case where $2k-3\leq|N|\leq 2k$. However, in the position $(x,2k,1)$ ($x\geq 2k+2$), the best move is instead to $(2k+1,2k,1)$, which is an element of $P^{+}\cap P^{-}$. For example, if the initial position is $(10,8,1)$ and $6\leq|N|\leq 8$, the game proceeds [^3] For $5\leq|N|\leq 6$, the game proceeds the same way up to $(10,8,1)\to(6,8,1)$, but then the second player makes a move to $(6,4,1)$. as follows:
@@
Remark 6.5.
While the payoff functions presented in Theorem 6.3 are symmetric with respect to $N=0$, payoff functions, in general, are not symmetric. In fact, as shown in Proposition 6.1, $f_{N}(x,x,z)$ is symmetric with respect to $N=-1$ rather than $N=0$, and $f_{N}(5,4,3)$ (as illustrated in Figure 6) lacks symmetry with respect to either $N=0$ or $N=-1$.
[…]
@@ Acknowledgements
This paper is a compilation of [4] and [5]. This resubmission is allowed under the guidelines of the Information Processing Society of Japan (For details, see https://www.ipsj.or.jp/english/faq/ronbun-faq_e.html). This work was supported by JSPS KAKENHI Grant Number JP21K12191.

### v2

@@ 6 Changes in the optimal strategy depending on the bonus value
Thus, it follows that $f_{N}(2m+1,2m,1)\geq 2$ for $N=-(2m-2),-(2m-4),\ldots,2m-4,2m-2$. Furthermore, since the slope of this function with respect to $N$ is at least 1-1 (as stated in Proposition 1(ii)), it follows that $f_{N}(2m+1,2m,1)\geq F_{m}(N)$ holds for all $N$.
Case 2: $N=\pm(2l+1)$ with $0\leq l<m-1$. We analyze the possible moves available to the first player at the position $(2m+1,2m,1)$ and explicitly describe the opponent’s responses. Through this analysis, we show that $f_{\pm(2l+1)}(2m+1,2m,1)$ does not exceed 1.
  • If the first player moves to $(2m+1,2m,0)$, then by (4.4)—with the second player responding by moving to (2m,2m,0)(2m,2m,0)—we obtain
    $1-f_{\pm(2l+1)}(2m+1,2m,0)=f_{\pm(2l+1)}(2m,2m,0)=1-|1+N|\leq 1.$
    Similarly, if the first player moves to $(2m,2m,1)$, the second player responds by moving to $(2m,2m,0)$.
  • If the first player moves to $(0,2m,1)$, the second player responds by moving to $(0,1,1)$, which is optimal, when $N=2l+1$. Then
    $2m+1-f_{2l+1}(0,2m,1)=2+f_{2l+1}(0,1,1)=2-N.$
    If N=(2l+1)N=-(2l+1), the second player responds by moving to $(0,0,1)$, which is optimal. Then
    $2m+1-f_{-(2l+1)}(0,2m,1)=1+f_{-(2l+1)}(0,0,1)=2+N.$
    Thus, in both cases, the function value does not exceed 1. Similarly, if the first player moves to $(1,2m,1)$, $(2m+1,0,1)$ or $(2m+1,1,1)$, the second player responds by moving to $(1,0,1)$ if $N=2l+1$, and to either $(1,1,1)$ or $(0,0,1)$ if N=(2l+1)N=-(2l+1), so that the function value does not exceed 1.
  • If the first player moves to $(x,2m,1)$ or $(2m+1,x,1)$ (where $x$ is an even number with $2\leq x\leq 2l$), the second player responds by moving to $(x,x+1,1)$ (in this sequence of moves, it is more advantageous to have initially moved to $(x,2m,1)$ for the first player). Then
@@
    By the induction hypothesis (a), we have $f_{\pm(2l+1)}(x,x+1,1)\leq-1$, implying that the right-hand side of the inequality does not exceed 1.
  • If the first player moves to $(x,2m,1)$ or $(2m+1,x,1)$ (where $x$ is an odd number with $3\leq x\leq 2m-1$), the second player responds by moving to (x,x1,1)(x,x-1,1) (it is more advantageous to have initially moved to $(x,2m,1)$ for the first player). Then
    $2m+1-x-f_{\pm(2l+1)}(x,2m,1)\leq f_{\pm(2l+1)}(x,x-1,1).$
    By the induction hypothesis (a), it follows that the right-hand side of the above inequality does not exceed 1.
  • If the first player moves to $(x,2m,1)$ or $(2m+1,x,1)$ (where $x$ is an even number with $2l+2\leq x\leq 2m-2$), the second player responds by moving to $(x,2l,1)$ (it is more advantageous to have initially moved to $(x,2m,1)$ for the first player). Then
@@
    By the induction hypothesis (b), it follows that the right-hand side of the above inequality does not exceed 1.
From the above moves and responses, we have established that $f_{N}(2m+1,2m,1)\leq 1$ for $N=-(2m-3),-(2m-5),\ldots,2m-5,2m-3$. Consequently, it follows that $f_{N}(2m+1,2m,1)\leq F_{m}(N)$ for $|N|\leq 2m-2$. On the other hand, by (5.1), it is known that for sufficiently large $|N|$, we have $f_{N}(2m+1,2m,1)=2m-|N|$. Thus, by Corollary 2, we conclude that for all $N$, $f_{N}(2m+1,2m,1)\leq 2m-|N|$. Hence, for all $N$, it follows that $f_{N}(2m+1,2m,1)\leq F_{m}(N)$.
Similarly, statements (ii) and (iii) for $k=m$ follow under the assumption that (i) holds for all $k\leq m$, and that (ii) and (iii) hold for all $k<m$. The details of the proof are given in the Appendix.  $\Box$
We provide additional insights into Theorem 11 and examine the factors contributing to the complexity of move selection and the increase in the number of breakpoints. First, as observed in the proof of (i), in the position $(2k+1,2k,1)$ ($k\geq 1$), the best move (or one of the best moves) for the player is to choose the even number $j$ which is closest to $N$ within the range $(2k-2)\leq j\leq 2k-2$ and move to $(|j|,2k,1)$. For example, in the position $(9,8,1)$, the player should make the following moves based on the value of $|N|$:
@@
This strategy is derived from (b), which states that $f_{N}(x,|j|,1)$ attains its minimum value x|j|1x-|j|-1 at $N=\pm|j|$. This represents the weakest positions for the current player. Thus, it is advantageous to force the opponent into such a position.
Additionally, in the positions $(2k+1,2k,1)$ and $(x,2k,1)$ ($x\geq 2k+2$), considered in (i) and (ii), a common strategy is observed when $|N|$ is small: the player takes stones from the largest pile so that the remaining number of stones is close to $|N|$. As a result, the function $F_{k}(N)$, with an upward or downward shift, appears in both payoff functions. The distinction between these two payoff functions lies in whether the player can force the opponent into a P-position when $|N|$ is large. Specifically, when $|N|\geq 2k$, in the position $(2k+1,2k,1)$, the optimal move is to (2k2,2k,1)(2k-2,2k,1), which is the same as in the case where $2k-3\leq|N|\leq 2k$. However, in the position $(x,2k,1)$ ($x\geq 2k+2$), the best move is instead to $(2k+1,2k,1)$, which is an element of $P^{+}\cap P^{-}$. For example, if the initial position is $(10,8,1)$ and $6\leq|N|\leq 8$, the game proceeds [^3] For $5\leq|N|\leq 6$, the game proceeds the same way up to $(10,8,1)\to(6,8,1)$, but then the second player makes a move to $(6,4,1)$. as follows:
@@
Remark 1.
While the payoff functions presented in Theorem 11 are symmetric with respect to $N=0$, payoff functions, in general, are not symmetric. In fact, as shown in Proposition 9, $f_{N}(x,x,z)$ is symmetric with respect to N=1N=-1 rather than $N=0$, and $f_{N}(5,4,3)$ (as illustrated in Fig. 6) lacks symmetry with respect to either $N=0$ or N=1N=-1.
0444-433NN
@@
Appendix A Proof of Theorem 11
In this Appendix, we provide a detailed proof of statements (ii) and (iii) of Theorem 11, whose proofs were omitted in Section 6. Specifically, we show that statements (ii) and (iii) for $k=m$ follow under the assumption that (i) holds for all $k\leq m$, and that (ii) and (iii) hold for all $k<m$.
To establish (ii), it suffices to consider integer values of $N$. It suffices to verify the following inequalities:
  • $f_{N}(x,2m,1)\geq x-2m+1$ when $N$ is even and $|N|\leq 2m-2$,
  • $f_{N}(x,2m,1)\geq x-4m-1+|N|$ when $|N|\geq 2m$,
  • $f_{N}(x,2m,1)\leq x-2m$ when $N$ is odd and $|N|\leq 2m-3$,
  • $f_{N}(x,2m,1)\leq x-2m-1$ when $N=\pm 2m$.
Case 1: $N=\pm 2l$ with $0\leq l<m$. The first player has the option to move from the position $(x,2m,1)$ to $(2l,2m,1)$. By the induction hypothesis (b) stated at the beginning of the proof of Theorem 11, we obtain
$f_{\pm 2l}(x,2m,1)\geq x-2l-f_{\pm 2l}(2l,2m,1)=x-2m+1.$
Case 2: $|N|\geq 2m$. The first player has the option to move from the position $(x,2m,1)$ to $(2m+1,2m,1)$. By the induction hypothesis, since (i) has already been established for $k=m$, we have $f_{N}(2m+1,2m,1)=2m-|N|$. Hence,
$f_{N}(x,2m,1)\geq x-(2m+1)-f_{N}(2m+1,2m,1)=x-4m-1+|N|.$
Case 3: $N=\pm(2l+1)$ with $0\leq l<m-1$. We analyze the possible moves available to the first player at the position $(x,2m,1)$ and explicitly describe the opponent’s responses. Through this analysis, we show that $f_{\pm(2l+1)}(x,2m,1)$ does not exceed x2mx-2m.
  • If the first player moves to $(2m,2m,1)$ and the second player responds by moving to $(2m,2m,0)$, then by (4.3) we obtain
    $x-2m-f_{\pm(2l+1)}(2m,2m,1)\leq x-2m-|1+N|\leq x-2m.$
    Similarly, if the first player moves to $(x,2m,0)$, the second player responds by moving to $(2m,2m,0)$.
  • If the first player moves to $(0,2m,1)$, the second player responds by moving to $(0,1,1)$ if $N=2l+1$. Then
    $x-f_{2l+1}(0,2m,1)\leq x-2m+1-N\leq x-2m$
    since $N$ is a positive odd integer. If N=(2l+1)N=-(2l+1), the second player responds by moving to $(0,0,1)$. Then
    $x-f_{-(2l+1)}(0,2m,1)\leq x-2m+1+N\leq x-2m$
    since $N$ is a negative odd integer. Similarly, if the first player moves to $(1,2m,1)$, $(2m+1,0,1)$ or $(2m+1,1,1)$, the second player responds by moving to $(1,0,1)$ if $N=2l+1$, and to either $(1,1,1)$ or $(0,0,1)$ if N=(2l+1)N=-(2l+1), so that the function value does not exceed x2mx-2m.
  • If the first player moves to $(x^{\prime},2m,1)$ or $(x,x^{\prime},1)$ (where $x^{\prime}$ is an even number with $2\leq x^{\prime}\leq 2l$), the second player responds by moving to $(x^{\prime},x^{\prime}+1,1)$ (in this sequence of moves, it is more advantageous to have initially moved to $(x^{\prime},2m,1)$ for the first player). Then
    $x-x^{\prime}-f_{\pm(2l+1)}(x^{\prime},2m,1)\leq x-2m+1+f_{\pm(2l+1)}(x^{\prime},x^{\prime}+1,1).$
    By the induction hypothesis (a), we have $f_{\pm(2l+1)}(x^{\prime},x^{\prime}+1,1)\leq-1$, implying that the right-hand side of the inequality does not exceed x2mx-2m.
  • If the first player moves to $(x^{\prime},2m,1)$ or $(x,x^{\prime},1)$ (where $x^{\prime}$ is an odd number with $3\leq x^{\prime}\leq 2m-1$), the second player responds by moving to $(x^{\prime},x^{\prime}-1,1)$ (it is more advantageous to have initially moved to $(x^{\prime},2m,1)$ for the first player). Then
    $x-x^{\prime}-f_{\pm(2l+1)}(x^{\prime},2m,1)\leq x-2m-1-f_{\pm(2l+1)}(x^{\prime},x^{\prime}-1,1).$
    By the induction hypothesis (a), it follows that the right-hand side of the above inequality does not exceed x2mx-2m.
  • If the first player moves to $(x^{\prime},2m,1)$ or $(x,x^{\prime},1)$ (where $2l+2\leq x^{\prime}$), the second player responds by moving to $(x^{\prime},2l,1)$ (it is more advantageous to have initially moved to $(x^{\prime},2m,1)$ for the first player). Then
    $x-x^{\prime}-f_{\pm(2l+1)}(x^{\prime},2m,1)\leq x-x^{\prime}-2m+2l+f_{\pm(2l+1)}(x^{\prime},2l,1).$
    By the induction hypothesis (b), it follows that the right-hand side of the above inequality does not exceed x2mx-2m.
Case 4: $N=\pm 2m$. A similar analysis of the available moves and responses yields $f_{\pm 2m}(x,2m,1)\leq x-2m-1$.
  • If the first player moves to $(x,2m,0)$ or $(2m,2m,1)$, the second player responds by moving to $(2m,2m,0)$.
  • If the first player moves to $(0,2m,1)$, the second player responds by moving to $(0,1,1)$, which is optimal, when $N=2m$. If N=2mN=-2m, the second player responds by moving to $(0,0,1)$, which is optimal. If the first player moves to $(1,2m,1),(2m+1,0,1)$ or $(2m+1,1,1)$, the second player responds by moving to $(1,0,1)$ if $N=2m$, and to either $(1,1,1)$ or $(0,0,1)$ if N=2mN=-2m. In each case, this response ensures that the function value does not exceed x2m1x-2m-1.
  • If the first player moves to $(x^{\prime},2m,1)$ or $(x,x^{\prime},1)$ (where $x^{\prime}$ is an even number with $2\leq x^{\prime}\leq 2m-2$), the second player responds by moving to $(x^{\prime},x^{\prime}+1,1)$ (in this sequence of moves, it is more advantageous to have initially moved to $(x^{\prime},2m,1)$ for the first player). Then
    $x-x^{\prime}-f_{\pm 2m}(x^{\prime},2m,1)\leq x-2m+1+f_{\pm 2m}(x^{\prime},x^{\prime}+1,1).$
    By the induction hypothesis (a), we have $f_{\pm 2m}(x^{\prime},x^{\prime}+1,1)\leq-2$, implying that the right-hand side of the inequality does not exceed x2m1x-2m-1.
  • If the first player moves to $(x^{\prime},2m,1)$ or $(x,x^{\prime},1)$ (where $x^{\prime}$ is an odd number with $3\leq x^{\prime}\leq 2m-1$), the second player responds by moving to $(x^{\prime},x^{\prime}-1,1)$ (it is more advantageous to have initially moved to $(x^{\prime},2m,1)$ for the first player). Then
    $x-x^{\prime}-f_{\pm 2m}(x^{\prime},2m,1)\leq x-2m-1-f_{\pm 2m}(x^{\prime},x^{\prime}-1,1).$
    By the induction hypothesis (a), it follows that the right-hand side of the above inequality does not exceed x2m1x-2m-1.
  • If the first player moves to $(x^{\prime},2m,1)$ (where $2m+2\leq x^{\prime}$), the second player responds by moving to $(2m+1,2m,1)$. Then
    $x-x^{\prime}-f_{\pm 2m}(x^{\prime},2m,1)\leq x-2x^{\prime}+2m+1+f_{\pm 2m}(2m+1,2m,1).$
    By the induction hypothesis, we have $f_{\pm 2m}(2m+1,2m,0)=0$, and since $2m+2\leq x^{\prime}$, the right-hand side of the above inequality does not exceed x2m1x-2m-1.
The proof of (iii) proceeds similarly. It suffices to verify the following inequalities:
  • $f_{N}(x,2m+1,1)\geq x-2m$ when $N$ is even and $|N|\leq 2m-2$,
  • $f_{N}(x,2m+1,1)\geq x-4m+|N|$ when $|N|\geq 2m$,
  • $f_{N}(x,2m+1,1)\leq x-2m-1$ when $N$ is odd and $|N|\leq 2m-1$.
Case 1: $N=\pm 2l$ with $0\leq l<m$. The first player has the option to move from the position $(x,2m+1,1)$ to $(2l,2m+1,1)$. By the induction hypothesis (b), we obtain
$f_{\pm 2l}(x,2m+1,1)\geq x-2l-f_{\pm 2l}(2l,2m+1,1)=x-2m.$
Case 2: $|N|\geq 2m$. The first player has the option to move from the position $(x,2m+1,1)$ to $(2m,2m+1,1)$. By the induction hypothesis, since (i) has already been established for $k=m$, we obtain
$f_{N}(x,2m+1,1)\geq x-2m-f_{N}(2m,2m+1,1)=x-4m+|N|.$
Case 3: $N=\pm(2l+1)$ with $0\leq l<m$. We analyze the possible moves available to the first player at the position $(x,2m+1,1)$ and explicitly describe the opponent’s responses. Through this analysis, we show that $f_{\pm(2l+1)}(x,2m+1,1)$ does not exceed x2m1x-2m-1. This case can be proved by the same procedure as in Case 3 of (ii). $\Box$
@@ Acknowledgements
This work was supported by JSPS KAKENHI (Grant Numbers JP21K12191 and JP25K15403) and JST SPRING (Grant Number JPMJSP2115).